# Embodied AI in Urban Spaces

### v1

@@ C.2 SocialNav Experiments
For SocialNav experiments, most settings are the same as the ones in PointNav. The most important difference is that dynamic agents will also be present in the environment. The trajectories of environmental agents are generated together by using the model of ORCA [77] with P&R [11]. Since vehicles are inherited from MetaDrive [44], we use the same parameter to control its density, i.e., traffic density $0.05$ in our experiments.
@@
For SocialNav, an episode is considered successful if the agent issues the DONE action, defined as completing 95% of the set route within 1,000 maximum steps. The agent is evaluated using the Success Rate (SR) and Social Navigation Score (SNS) [12], which is the average of Success weighted by Time Length (STL) and Personal Space Compliance (PSC). SNS measures the agent in terms of safety and efficiency.
@@ Methods.
Appendix D Datasheet
Table 8: A datasheet [25] for MetaUrban and MetaUrban-12K.
@@ Appendix D Datasheet
For what purpose was the dataset created? The dataset was created to enable agents training on diverse scenes and facilitate Embodied AI research in urban spaces.
@@
Over what timeframe was the data collected? Data was collected in May 2024.
@@
What (other) tasks could the dataset be used for? The scenes can be used in a wide variety of tasks in embodied AI, computer vision, and urban planning.
@@
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? For 3D human assets, we use Synbody [81]. Its license is CC BY-NC-SA 4.0. For movement sequences, we use BEDLAM [6]. See https://bedlam.is.tue.mpg.de/license.html for its license.
@@ Appendix E Performance
Appendix F Discussion
@@
Limitations. 1) Real-world scene distribution. In this work, we extract object category distribution from real-world data of urban spaces. Other than the real-world distribution of object categories, the distribution of object location and scene layout is also important for constructing specialized scenes for agent training. Extraction of such distribution relies on an accurate reconstruction of 3D scenes from real-world videos or even images, and thus is extremely challenging. An interesting direction is extracting real-world scene distribution from in-the-wild videos, including object category, object location, and scene layout. Then, we can build a digital twin of a target scene for the agent’s training. It could help to develop scene-specific agents. 2) Interactive agent behaviors. In this work, we construct the environmental agents‘ dynamic with deterministic methods, determining their movements and trajectories with rules. However, in the real world, all environmental agents are interactive; their behaviors are affected by each other and the surrounding environments. An interesting research direction is to endow personal traits like job, personality, and purpose to agents and harness the advances of LLMs [1] and LVMs [45] to form social [62] and interactive behaviors [55] of agents in urban scenes spontaneously. 3) Robot-specific tasks. In this work, we design point navigation and social navigation tasks as the pilot study using MetaUrban for embodied AI research. However, different robots (delivery bots and humanoid robots) have significantly different kinematics and functions. Thus, more tasks, such as locomotion and manipulation, can be explored. An important direction is to design different tasks based on different robots, which could enable various complex but crucial services in urban environments. 4) Efficiency. In this work, different from indoor scenes and driving simulators, MetaUrban supports generating complex interactive urban scenes with arbitrary scales. However, with the increase in scale, the number of objects and dynamic agents will surge dramatically, which will bring the degradation of the efficiency of physical simulation and rendering. A promising direction is to integrate more sophisticated physical engines and renders, such as NVIDIA Omniverse.
@@
Future work. 1) Foundation model. MetaUrban can easily generate infinite urban scenes with a large quantity of semantics and complex interactions, which could facilitate the pre-training of foundation models (like LLMs and LVMs) that can be used for downstream agent learning tasks. 2) Human-robot cohabitate. Mobile machines have started emerging in the urban space, which makes it no longer exclusive to humans. We plan to work with urban sociologists to study the influence of robots on human urban life through both simulation and field experiments. 3) Improve limitations. The directions discussed in limitations – real-world scene distribution extraction, interactive agent behavior simulation, robot-specific task design, and efficiency optimization are also meaningful directions. In summary, MetaUrban, as a new urban environment simulator, will bring a lot of new interesting research directions. We are dedicated to maintaining MetaUrban in the long term and supporting the community’s efforts to develop it into a sustainable infrastructure.

### v2

@@ D.2 SocialNav Experiments
For SocialNav experiments, most settings are the same as the ones in PointNav. The most important difference is that dynamic agents will also be present in the environment. The trajectories of environmental agents are generated together by using the model of ORCA (Van Den Berg et al., 2011) with P&R (De Wilde et al., 2014). Since vehicles are inherited from MetaDrive (Li et al., 2022b), we use the same parameter to control its density, i.e., traffic density $0.05$ in our experiments.
@@
For SocialNav, an episode is considered successful if the agent issues the DONE action, defined as completing 95% of the set route within 1,000 maximum steps. The agent is evaluated using the Success Rate (SR) and Social Navigation Score (SNS) (Deitke et al., 2022a), which is the average of Success weighted by Time Length (STL) and Personal Space Compliance (PSC). SNS measures the agent in terms of safety and efficiency.
@@ Methods.
D.3 Evaluation across Mobile Machines
In this experiment, we evaluate the influence of mechanical structures in policy execution on different terrains. We conduct experiments on three types of wheeled mobile machines – a delivery bot, an electric wheelchair, and a mobility scooter, with remarkably different specifications, such as wheelbase, wheel radius, and wheel width. We designed three sufficiently long runways with three kinds of terrains: slopes, stairs, and roughs. From the starting point, each runway has a gradually increasing difficulty. Slopes will have an increasingly steeper angle; stairs will have increasingly higher step heights; roughs will have increasingly larger bumps. We apply the same "moving forward" policy to each mobile machine to test the longest distance and time duration they can travel before termination, and report the metrics of $x$-displacement (m) and Time to fall (s) (Agarwal et al., 2023) respectively. Terminal conditions are getting stuck, slowing down significantly, and toppling over.
Results are shown in Table 9. The mobility scooter, which has the largest wheelbase, wheel width, and radius, achieves the best performance in the slopes test. It indicates that a larger wheelbase increases stability and reduces the risk of tipping backward on steep inclines, while wheel width and radius help in better traction on slopes. All three machines show similar but poor performance in the stairs test. It indicates the inherent defect of wheeled mobile machines and emphasizes the importance of accessibility in public urban spaces. The delivery bot, which has the smallest wheelbase, wheel width, and radius, achieves poor performance on all three tests. It indicates that although the delivery bot’s structure gives it good maneuverability on flat surfaces, it comes at the cost of losing stability on complex terrains.
Table 9: Evaluation of policy execution across mobile machines. For each row of different terrains,   indicates the best performance among the three machines.
Terrain $x$-displacement (m) $\uparrow$ Time to fall (s) $\uparrow$
Delivery Bot Wheelchair Mobility Scooter Delivery Bot Wheelchair Mobility Scooter
Wheelbase ($m$) 0.45 0.5 0.6 0.45 0.5 0.6
Wheel radius ($m$) 0.1 0.15 0.2 0.1 0.15 0.2
Wheel width ($m$) 0.1 0.1 0.15 0.1 0.1 0.15
Slopes 31.90 34.58 38.07 6.3 6.9 7.7
Stairs 38.55 38.94 38.67 7.0 7.8 7.2
Roughs 28.06 31.91 34.17 5.8 6.4 6.7
Appendix E Unique Challenges in Urban Micromobility
In this section, we delve into and validate four unique challenges a mobile machine will encounter in public urban space, which is the stage of urban micromobility tasks, distinct from previous indoor and driving environments, i.e., long horizon tasks in large-scale scenes, multifarious terrains, diverse obstacles, and dense pedestrians.
[…]
Table 10: Unique challenges validation. (Left) Long horizon tasks in large-scale scenes. (Right) Dense pedestrians.
Setting-1 Setting-2
[…]
Setting-1 Setting-2 Setting-3 Setting-4
[…]
Multifarious terrains. In indoor environments, most of the grounds are smooth and even, while in driving environments, roadway surfaces can only have slight cracks and damage. Yet, in public urban spaces, mobile machines will encounter multifarious complex terrains, such as slopes, stairs, and roughs. As shown in Table 9, different terrains will deeply influence the performance of mobile machines with different mechanical structures, such as wheelbase, wheel radius, and width, etc.. Results show that along with the increasing difficulty of terrains, all of the machines will fail because of getting stuck, barely moving, or toppling over. It indicates that multifarious terrains bring a unique challenge to mobile machines, and validate the necessity of the terrain generation system designed in MetaUrban.
Diverse obstacles. In indoor environments, even though there are many objects, the distribution has a large variation compared to urban spaces. In driving environments, there are only a few obstacles on roadways, such as cones and barriers. As shown in Figure 26, we compare the distribution of object category in MetaUrban with that in ProcThor (Deitke et al., 2022b), a state-of-the-art indoor simulation environment. We can observe a significant variation in object category distributions between public urban spaces and indoor spaces, although these two scenarios both accommodate a lot of objects. The diversity, particularity, and concentration of obstacles in urban spaces present a unique challenge for mobile machines. The statistical results validate the necessity of the pipeline of scalable obstacle filling in MetaUrban.
Figure 26: Object category distributions. Red: Distribution of object category in urban spaces. Blue: Distribution of object category in indoor spaces.
Dense pedestrians. In indoor environments, humans will share walkable spaces with humans; however, there are only 3-5 people in one room in common (as shown in Habitat 3.0 (Puig et al., 2023b)). In driving environments, except the intersections, there exist barely any shared spaces for pedestrians and vehicles. In contrast, in public urban spaces, almost all of the spaces for mobile machines are shared with pedestrians. We compare models trained with different pedestrian densities and locations. Setting-1: 10 pedestrians per 100-meter episode (a common scenario in an urban environment). Setting-2: 5 pedestrians per 100-meter episode (a common scenario in an indoor environment). Setting-3: 10 pedestrians per 100-meter episode but only shown in intersections (a common scenario in a driving environment). All the three models are then tested on the MetaUrban-test dataset. Setting 4: using the model trained in Setting-1 and testing on a density of 30 pedestrians per 100-meter episode (a crowd scenario in an urban environment).
Results are shown in Table 10 (Right). We take Setting-1 as a reference. With fewer pedestrians in Setting-2, the success rate will decrease, and the cost will increase significantly, indicating a higher frequency of bumping into pedestrians. With different distribution but the same pedestrian density in Setting-3, both success rate and cost degrade dramatically, indicating the unique challenge of sharing walkable regions with pedestrians. In Setting 4, we further increase the pedestrian’s density and see a huge degradation in success rate but a moderate degradation in Cost. It indicates that having more pedestrians poses a significant challenge for the agent to reach the goal point. Interestingly, the agent still attempts to avoid pedestrians due to its effective training in public urban spaces. Results demonstrate that the high pedestrian density and interaction frequency in public urban spaces will place a unique challenge for mobile machines. It also validates the importance of the Cohabitant Populating module in MetaUrban.
Appendix F Datasheet
Table 11: A datasheet (Gebru et al., 2021) for MetaUrban and MetaUrban-12K.
@@ Appendix F Datasheet
For what purpose was the dataset created? The dataset was created to enable agents training on diverse scenes and facilitate AI-driven urban micromobility research.
@@
Over what timeframe was the data collected? Data was collected in Sept. 2024.
@@
What (other) tasks could the dataset be used for? The scenes can be used in a wide variety of tasks in urban micromobility, embodied AI, computer vision, and urban planning.
@@
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? For 3D human assets, we use Synbody (Yang et al., 2023). Its license is CC BY-NC-SA 4.0. For movement sequences, we use BEDLAM (Black et al., 2023). See https://bedlam.is.tue.mpg.de/license.html for its license.
@@ Appendix G Performance
Appendix H Robustness
We trained PPO on PointNav with different seeds and found that the variance of the performance across different seeds is small. The success rate of the PPO agents is $0.695\pm 0.014$ on the MetaUrban-test set and $0.638\pm 0.060$ on the MetaUrban-unseen set.
Appendix I Discussion
@@
Limitations. 1) Real-world scene distribution. In this work, we extract object category distribution from real-world data of urban spaces. Other than the real-world distribution of object categories, the distribution of object location and scene layout is also important for constructing specialized scenes for agent training. Extraction of such distribution relies on an accurate reconstruction of 3D scenes from real-world videos or even images, and thus is extremely challenging. An interesting direction is extracting real-world scene distribution from in-the-wild videos, including object category, object location, and scene layout. Then, we can build a digital twin of a target scene for the agent’s training. It could help to develop scene-specific agents. 2) Interactive agent behaviors. In this work, we construct the environmental agents‘ dynamic with deterministic methods, determining their movements and trajectories with rules. However, in the real world, all environmental agents are interactive; their behaviors are affected by each other and the surrounding environments. An interesting research direction is to endow personal traits like job, personality, and purpose to agents and harness the advances of LLMs (Achiam et al., 2023) and LVMs (Liu et al., 2023a) to form social (Puig et al., 2023a) and interactive behaviors (Park et al., 2023) of agents in urban scenes spontaneously. 3) Robots’ additional capability learning. In urban micromobility, safe navigation through the city is the primary goal for mobile machines. However, additional capabilities, such as locomotion and manipulation, can enable robots to perform more complex tasks in urban spaces. Thus, an important direction is to extend MetaUrban to support additional capabilities learning gradually. It could enable various complex but important services in urban environments. 4) Efficiency. In this work, different from indoor scenes and driving simulators, MetaUrban supports generating complex interactive urban scenes with arbitrary scales. However, with the increase in scale, the number of objects and dynamic agents will surge dramatically, which will bring the degradation of the efficiency of physical simulation and rendering. A promising direction is to integrate more sophisticated physical engines and renders.
Sim-to-Real. MetaUrban is an embodied AI simulator that can enable fast iterative model development and evaluation before deployment. However, as it is impossible to perfectly replicate the real world without some loss, sim-to-real gap is an inevitable issue faced by all embodied AI simulators (Kolve et al., 2017; Savva et al., 2019; Shen et al., 2021; Li et al., 2024; Szot et al., 2021; Deitke et al., 2020; 2022b; Krajzewicz et al., 2002; Li et al., 2022b; Dosovitskiy et al., 2017), which also exists in MetaUrban. Following the standard setting embodied AI simulators, i.e., AI2-THOR (Kolve et al., 2017), Habitat (Savva et al., 2019), and ProcTHOR (Deitke et al., 2022b), we did not focus on real-world experiments so far.
However, positioning MetaUrban as a long-term maintained and sustainable community infrastructure, we do think the sim-to-real gap is a fundamental issue for simulators and are continuously working in two directions to tackle this problem. 1) Building real-world evaluation platform. Exceeding from an embodied AI platform, we are constructing an end-to-end experimental platform that spans from the MetaUrban simulation to the real-world deployment of robots, to progressively improve the designs in the simulator. We have conducted experiments in Unitree’s Go2 quadruped robot and COCO Robotics’s wheeled robot. We found that using more abstract observations (such as semantic maps and LiDAR maps provided in MetaUrban) to train robots with domain randomization can already achieve good generalizability to real-world environments. 2) Improving the realism of the simulation environment. First, to improve appearance realism, we are working on integrating Nvidia OmniVerse as the renderer. Then, to improve physical realism, we are working on integrating Nvidia PhysX into MetaUrban.
@@
Future work. 1) Foundation model. MetaUrban can easily generate infinite urban scenes with a large quantity of semantics and complex interactions, which could facilitate the pre-training of foundation models (like LLMs and LVMs) that can be used for downstream agent learning tasks. 2) Human-robot cohabitate. Mobile machines have started emerging in the urban space, which makes it no longer exclusive to humans. We plan to work with urban sociologists to study the influence of robots on human urban life through both simulation and field experiments. 3) Improve limitations. The directions discussed in limitations and sim-to-real gaps are also meaningful future work we will conduct. In summary, MetaUrban, as a new urban environment simulator, will bring a lot of new interesting research directions. We are dedicated to maintaining MetaUrban in the long term and supporting the community’s efforts to develop it into a sustainable infrastructure.